OpenStudy (anonymous):

Check my answer? What expression can be written as 4 • (3 + 8)? My answer is 4 • 3 + 8.

hartnn (hartnn):

nopes u distributed incorrectly

hartnn (hartnn):

a(b+c)= ab+ac

hartnn (hartnn):

so 4(3+8) = 4*3+4*8 ok ?